A circle has diameter 25 cm. how far from the centre of this circle is a chord 16 cm long?
Answer by mananth(16946)

draw a perpendicular from the center to the chord.
the line bisects the chord.
1/2 chord = 8 cm
the radius the segment of chord and the perpendicular form a right triangle
let distance be x
x^2+8^2= 25^2 ( By Pythagoras theorem)
x^2= 25^2-8^2
x^2=625-64
x^2= 561
x= 23.7 cm